***I need the steps to solve this problem by Sunday*** Please HELP!!!
The answer to the problem is: S(h)=6h-50
The problem is:
SKYLINE ORDINANCE A city council is proposing a new skyline ordinance. It would require the setback (S) for any building from a residence to be a minimum of 100 feet, plus an additional 6 feet for each foot of height above 25 feet. Find a linear function for (S) in terms of (h).
Thank you!

Answer by stanbon(75887)

The answer to the problem is: S(h)=6h-50
The problem is:
SKYLINE ORDINANCE A city council is proposing a new skyline ordinance. It would require the setback (S) for any building from a residence to be a minimum of 100 feet, plus an additional 6 feet for each foot of height above 25 feet. Find a linear function for (S) in terms of (h).
----
If height is 25, S = 100
If height is 26, S = 106
---
Figure the slope:
m = (106-100)/(26-25) = 6
---
S = mh + b
100 = 6*25 + b
100 = 150 + b
b = -50
----
Equation:
S(h) = 6h-50 for h a whole number >= 25.
=========================================
Cheers,
Stan H.
==============